howing yor cakcalations, and writein cnpo Now convert your answer into magnitude and drection. Record your answer here F,-( 一一, N , angle of Question: Are your answers from the graphical l and the algebraic methed consistent wih each other? Which of the two metheds do you think is more axewane? Explain


continued from previous question
Answer #1

Here, F1 = (20 i + 30 j) N

           F2 = (14 i + 25 j) N

=>   F3 = F1 + F2

               =   (20 i + 30 j) + (14 i + 25 j)

                = ( 34 i + 55 j) N

=> magnitude of F3 = sqrt(342 + 552)

   =      64.66 N

=>   direction   = tan-1(55/34)

                        = 58.27 degree